### v1

# Fermi Velocity Dependent Critical Current in Ballistic Bilayer Graphene Josephson Junctions

Abstract
We perform transport measurements on proximitized, ballistic, bilayer graphene Josephson junctions (BGJJs) in the intermediate-to-long junction regime ($L>\xi$). We measure the device’s differential resistance as a function of bias current and gate voltage for a range of different temperatures. The extracted critical current $I_{C}$ follows an exponential trend with temperature: $\exp(-k_{B}T/\delta E)$. Here $\delta E=\hbar\nu_{F}/2\pi L$: an expected trend for intermediate-to-long junctions. From $\delta E$, we determine the Fermi velocity of the bilayer graphene, which is found to increase with gate voltage. Simultaneously, we show the carrier density dependence of $\delta E$, which is attributed to the quadratic dispersion of bilayer graphene. This is in contrast to single layer graphene Josephson junctions, where $\delta E$ and the Fermi velocity are independent of the carrier density. The carrier density dependence in BGJJs allows for additional tuning parameters in graphene-based Josephson Junction devices.
Ballistic graphene Josephson junctions (GJJs) have been widely utilized as a platform to study various interesting physics that arise at low temperatures from interactions between superconductors and normal metals [1, 2, 3, 4, 5, 6]. The ballistic superconductor-normal metal-superconductor Josephson junction (SNSJJ) hosts Andreev bound states (ABS) which carry supercurrents across the normal region of the JJ; a disorder-free weak link and high transparency at the SN interface are necessary. Hexagonal Boron-Nitride (hBN) encapsulated graphene as the weak link enables highly transparent contacts at the interface whilst keeping graphene clean throughout the fabrication process [7]. Here, we study proximitized, ballistic, bilayer graphene Josephson junctions (BGJJs). Bilayer graphene devices (in contrast to monolayer) allow extra potential tunability via a non-linear dispersion relation, applied displacement field, or lattice rotation [6].
The critical current ($I_{C}$) of SNSJJ in the intermediate-to-long regime, where the junction length (L) $\geq$ superconducting coherence length ($\xi_{0}$), scales with temperature (T) as $I_{C}=exp(-k_{B}T/\delta E)$. Here, $\delta E=\hbar\nu_{F}/2\pi L$, an energy scale related to ABS level spacing [8, 9, 10, 11, 12]. Note that in the intermediate regime ($L\approx\xi_{0}$) $\delta E$ is found to be suppressed [4]. A previous study of GJJs found that in this regime, the relation is held more precisely when $\xi$ was taken into account along with L, that is: $\delta E=\hbar\nu_{F}/2\pi(L+\xi)$ [8]. Monolayer graphene displays a linear dispersion relation, which results in a constant fermi velocity ($\nu_{F0}$). Thus, in ballistic GJJs, $\delta E$ remains independent of the carrier density. In comparison, bilayer graphene displays a quadratic dispersion relation at low energies. In BGJJs we studied, a back-gate voltage ($V_{G}$) controls the carrier density; and $\delta E$ dependence on $V_{G}$ is observed. Using $\delta E$, we extract Fermi velocity in bilayer graphene: it is seen that $\nu_{F}$ increases with $V_{G}$, and saturates to the constant value, $\nu_{F0}$, of the monolayer graphene.
Our device consists of a series of four terminal Josephson junctions (on $\mathrm{S}iO_{2}/Si$ substrate) made with hBN encapsulated bilayer graphene contacted by Molybdenum-Rhenium (MoRe) electrodes. Bilayer graphene is obtained via the standard exfoliation method. It is then encapsulated in hexagonal Boron-Nitride using the dry transfer method [13]. MoRe of $80$ nm thickness is deposited via DC magnetron sputtering. The resulting device has four junctions of lengths $400$ nm, $500$ nm, $600$ nm, and $700$ nm. The width of the junctions is $4~{}\mu$m. The device is cooled in a Leiden cryogenics dilution refrigerator operated at temperatures above $1$ K, and the measurements were performed using the standard four-probe lock-in method. A gate voltage $V_{G}$ is applied to the $Si$ substrate with the oxide layer acting as a dielectric, which allows modulation of the carrier density.[14, 15, 4, 8, 16, 17].
Figure 1: (a) Differential resistance versus gate voltage ($V_{G}$) and bias current $I_{Bias}$ taken at $T=1.37$ K. The black region around zero bias corresponds to the superconducting state. $I_{Bias}$ is swept from negative to positive. Thus, the transition at negative bias corresponds to the re-trapping current $I_{R}$, while the transition at positive bias is the switching current $I_{C}$. (b) Vertical line cut of the resistance map taken at $V_{G}=5$ V, $T=1.37$ K, showing device’s differential resistance versus bias current.
Fig. 1(a) displays the differential resistance ($dV/dI$) map of the $400$ nm junction at $T=1.37$ K; we see zero resistance (black region) across all applied $V_{G}$ indicating the presence of supercurrent. As the bias current $I_{bias}$ is swept from negative to positive values, the junction first reaches its superconducting state at a value $|I_{bias}|=I_{R}$, known as re-trapping current. Then, as $|I_{bias}|$ is increased to higher positive values, the junction transitions to the normal state at $|I_{bias}|=I_{S}$, known as switching current. Fig. 1(a) shows that the junction can sustain larger region of critical current as we modulate the carrier density to higher values via $V_{G}$. Fig. 1(b) displays a line plot extracted from the $dV/dI$ map which shows hysteresis in $I_{R}$ and $I_{S}$. This is a commonly observed phenomenon in underdamped junctions [18, 14], or can also be attributed self-heating [15, 19, 16]. The critical current $I_{C}$ of the junction is approximately $I_{S}$ as found in switching statistics measurements, discussed in previous publications [8, 20, 21, 22].
Extracting the critical current $I_{C}$ from the differential maps for different temperatures, we can see that $I_{C}$ falls exponentially with inverse $T$ (Fig. 2c) We also extract conductance of the junction in the normal regime ($I_{Bias}\gg I_{C}$). Fig. 2(b) shows this conductance ($G$) for $400$ nm junction device. We find that conductance $G$ scales as the square-root of $V_{G}$ as seen from the fit (blue curve). However, due to large contact resistance ($R_{C}$) of the device, the measured conductance $G$ is suppressed compared to the ballistic limit expectation. Therefore, to demonstrate the ballistic nature of the device, we present normal resistances ($R_{N}$) of junctions of length $500$ nm, $600$ nm, and $700$ nm (Fig. 2(c) inset). Taking into account $R_{C}$, we plot extracted values versus $V_{G}$ from the Dirac point. The inset plot shows that the $R_{N}-R_{C}$ values are independent of junction length, which demonstrates the ballistic nature of the devices.
Figure 2: (a) Device picture. Image shows series of junctions with different Lengths: $400$ nm, $500$ nm, $600$ nm and $700$ nm. (b) The ballistic conductance vs Gate voltage for $L=400$ nm junction. The inset shows junction resistance minus the parasitic contact resistance plotted against gate voltage from the Dirac point for all our devices. (c) Critical currents $I_{C}$ of $L=400$ nm junction plotted against temperature $T$, for various gate voltages, on a semi-log scale. The plots show $V_{G}$ dependence of $I_{C}$: the gray lines show that the slope of the curve for the lowest plotted gate $V_{G}=8$ V, is smaller than the slope of the highest plotted gate $V_{G}=21$ V.
To extract $\delta E$ of the junction, we go to the discussion of $I_{C}$ vs temperature trends in Fig. 2(c). Here the y-axis is plotted in log scale. From the slope of curves $\mathrm{Log}(I_{C})=-(k_{B}/\delta E)T$ for each gate, one can extract $\delta E$ versus $V_{G}$ (plotted in Fig. 3a). Unlike for the case of monolayer graphene, a clear dependence on $V_{G}$ is seen. The energy $\delta E$ scales linearly with Fermi Velocity $v_{F}$ (Fig. 3b). Note that calculating $v_{F}$ from $\delta E$ for junctions in the intermediate regime requires knowledge of the superconducting coherence length $\xi$. In our case $\xi$ is obtained from the fit discussed below.
We now compare the experimentally obtained $\delta E$ (and $v_{F}$) to the theoretical expectation. Assuming a quadratic dispersion relation $\mathcal{E}_{F}=\frac{\hbar^{2}k^{2}_{F}}{2m^{*}}$, it follows that the expression for the Fermi velocity is: $v_{F}=\sqrt{\frac{2E_{F}}{m^{*}}}$ [23, 24, 25]. The Fermi Energy $\mathcal{E}_{F}$ for bilayer graphene scales as: $\mathcal{E}_{F}=\frac{\hbar^{2}\pi|n|}{2m^{*}}$. With $m^{*}$ being the effective mass of electrons in graphene. The carrier concentration $n$, controlled by the applied gate voltage $V_{G}$, is given by $n=\frac{V_{G}-V_{d}}{e}C_{Total}$ with $V_{d}$ as the gate voltage at the Dirac point. The total capacitance $C_{Total}$ is a combination of quantum capacitance $C_{q}$ and gate oxide capacitance $C_{ox}$: $C_{Total}=\left[\frac{1}{C_{ox}}+\frac{1}{C{q}}\right]^{-1}$. The quantum capacitance $C_{q}$ for bilayer graphene is determined by $C_{q}=\frac{2e^{2}m^{*}}{\pi\hbar^{2}}$. The gate oxide capacitance per unit area is $C_{ox}=\frac{\epsilon_{0}\epsilon_{r}}{d}$, where $\epsilon_{0}$ is vacuum permittivity, $\epsilon_{r}$ is the relative permittivity of the oxide, and $d$ is oxide layer thickness. For Silicon oxide gate with $d=300$ nm we get $C_{ox}\approx 115\mu F/m^{2}$. Thus, the full expression for the Fermi velocity $v_{F}$ is:
$v_{F}=\hbar\sqrt{\frac{2\pi\epsilon_{0}\epsilon_{r}e|V_{G}-V_{d}|}{m^{*}(2de^{2}m^{*}+\pi\epsilon_{0}\epsilon_{r}\hbar^{2})}}$ (1)
Note that the effective mass $m*$ typically ranges from $0.024~{}m_{e}$ to $0.058~{}m_{e}$ for $1*10^{12}\sim 4*10^{12}$ carriers/$cm^{2}$ [26], where $m_{e}$ is the electron rest mass. Moreover, since $m*$ has a carrier concentration dependence, we assume a linear shift of $m*$ with the gate voltage as: $m^{*}=m_{i}+dm(V_{G}-V_{d})$ [26].
Experimental data provides us with the following: $\delta E(V_{G})=\frac{\hbar}{2\pi(L+\xi)}v_{F}$. To fit this data, the model is set as :$\delta E(V_{G})=\mathcal{F}(m_{i},dm,\xi,V_{d},d)$ where $m_{i},dm,\xi,V_{d},d$ are the fitting parameters, and $V_{G}$ is the independent variable. (We use the as-designed length of the device $L$, and take $\epsilon_{r}=3.9$ for $SiO_{2}$.)
The resulting fit of the data from the $400~{}nm$ junction for both $\delta E$ and $v_{F}$ is plotted as solid lines in Fig. 3(a) and Fig. 3(b) respectively. Moreover, taking the fitted $\xi$, we calculate the Fermi velocity $v_{F}$ for all other junctions on the same substrate. As seen from Fig. 3(b), the calculated $v_{F}$ of all devices is in good agreement with the fit obtained from the $400$ nm junction. The fitted parameters are summarized in Table 1. All fall within the range of expected values, with $\xi$ being consistent with previously measured values for graphene/MoRe junctions.
Parameter Fitted value Expected Value
$\xi$ $460.34$ nm $300\sim 500$ nm
$d$ $317.71$ nm $300\sim 330$ nm
$m_{i}$ $0.020~{}m_{e}$ $0.02\sim 0.06~{}m_{e}$
$dm$ $0.0003~{}m_{e}$/V $0.0002\sim 0.0005~{}m_{e}$/V
$V_{d}$ $-1.99$ V ±$\approx\pm 2$ V
Table 1: The fitting parameters used to match the measured $\delta E$, and consequently the Fermi velocity $v_{F}$, versus gate to the theoretical expectation described in Equation 1. We see that resulting fitted values match closely to what is expected. The expected gate dielectric thickness $d$ is estimated from the substrate specifications plus the bottom hBN thickness. The expected Dirac point voltage $V_{D}$ is ontained from the resistance map. The expectations for the superconducting coherence length $\xi$ and the effective mass $m_{i}$ are obtained from previous works[26, 4].
Figure 3: (a) Energy $\delta E$ extracted from the slope of log($I_{C}$) vs T plotted against the gate voltage $V_{G}$ from the Dirac point of the junction with $L=400$ nm. We see $\delta E$ dependence on the carrier density modulated via the gate voltage for the junction. (b) Fermi velocity ($v_{f}$) calculated from $\delta E$ using the device dimensions, and the superconducting coherence length $\xi$ obtained from the fit to theory. The solid line represents the theoretical trend as fitted to the data for the $L=400$ nm junction. In addition, panel (b) shows calculated $v_{F}$ for the other junctions using the $\xi$ from the $L=400$ nm fit.
In conclusion, we study the evolution of critical current with respect to gate in bilayer graphene Josephson Junctions (BGJJs). Using the critical current-temperature relation expected for intermediate-to-long junctions, we extract the relevant energy scale $\delta E$ and find that it has a clear gate dependence. As $\delta E$ is proportional to the Fermi velocity $v_{F}$ in the bilayer graphene, we are able to match the observed gate dependence to the theoretical expectation. Our observation is contrasted with monolayer graphene JJs, that do not have a gate dependent $\delta E$. This result showcases the greater tunability of BGJJs, and offers additional avenues for device characterization.

### v2

@@
 The carrier density dependence in BGJJs allows for additional tuning parameters in graphene-based Josephson Junction devices.
-Ballistic graphene Josephson junctions (GJJs) have been widely utilized as a platform to study various interesting physics that arise at low temperatures from interactions between superconductors and normal metals [1, 2, 3, 4, 5, 6].
-The ballistic superconductor-normal metal-superconductor Josephson junction (SNSJJ) hosts Andreev bound states (ABS) which carry supercurrents across the normal region of the JJ; a disorder-free weak link and high transparency at the SN interface are necessary.
-Hexagonal Boron-Nitride (hBN) encapsulated graphene as the weak link enables highly transparent contacts at the interface whilst keeping graphene clean throughout the fabrication process [7].
+keywords:
+Graphene, Bilayer Graphene, Josephson Junctions, Fermi Velocity, Andreev Levels.
+Ballistic graphene Josephson junctions (GJJs) have been widely utilized as a platform to study novel quantum physics phenomena1, 2 and devices3, including: entangled pair generation4, 5, topological states arising from the mixing of superconductivity and quantum Hall states6, as well as photon sensing via bolometry/calorimetry7.
+Superconductor-normal metal-superconductor Josephson junction (SNSJJ) hosts Andreev bound states (ABS), which carry supercurrents across the normal region of the JJ; in order to enter the ballistic regime, a disorder-free weak link and high transparency at the SN interface are necessary.
+Hexagonal Boron-Nitride (hBN) encapsulated graphene as a weak link enables highly transparent contacts at the interface whilst keeping graphene clean throughout the fabrication process 8.
 Here, we study proximitized, ballistic, bilayer graphene Josephson junctions (BGJJs).
-Bilayer graphene devices (in contrast to monolayer) allow extra potential tunability via a non-linear dispersion relation, applied displacement field, or lattice rotation [6].
+Bilayer graphene devices (in contrast to monolayer) allow extra potential tunability via a non-linear dispersion relation, applied displacement field, or lattice rotation 1.
 The critical current ($I_{C}$) of SNSJJ in the intermediate-to-long regime, where the junction length (L) $\geq$ superconducting coherence length ($\xi_{0}$), scales with temperature (T) as $I_{C}=exp(-k_{B}T/\delta E)$.
-Here, $\delta E=\hbar\nu_{F}/2\pi L$, an energy scale related to ABS level spacing [8, 9, 10, 11, 12].
-Note that in the intermediate regime ($L\approx\xi_{0}$) $\delta E$ is found to be suppressed [4].
-A previous study of GJJs found that in this regime, the relation is held more precisely when $\xi$ was taken into account along with L, that is: $\delta E=\hbar\nu_{F}/2\pi(L+\xi)$ [8].
+Here, $\delta E=\hbar\nu_{F}/2\pi L$, an energy scale related to the ABS level spacing 2, 9, 10, 11, 12, 13.
+Note that in the intermediate regime ($L\approx\xi_{0}$) $\delta E$ is found to be suppressed 5.
+A previous study of GJJs found that in this regime, the relation is held more precisely when $\xi$ was taken into account along with L, that is: $\delta E=\hbar\nu_{F}/2\pi(L+\xi)$ 2, 13.
 Monolayer graphene displays a linear dispersion relation, which results in a constant fermi velocity ($\nu_{F0}$).
@@
 In BGJJs we studied, a back-gate voltage ($V_{G}$) controls the carrier density; and $\delta E$ dependence on $V_{G}$ is observed.
-Using $\delta E$, we extract Fermi velocity in bilayer graphene: it is seen that $\nu_{F}$ increases with $V_{G}$, and saturates to the constant value, $\nu_{F0}$, of the monolayer graphene.
+Using $\delta E$, we extract the Fermi velocity in bilayer graphene: It is seen that $\nu_{F}$ increases with $V_{G}$, and saturates to the constant value, $\nu_{F0}$, of the monolayer graphene.
 Our device consists of a series of four terminal Josephson junctions (on $\mathrm{S}iO_{2}/Si$ substrate) made with hBN encapsulated bilayer graphene contacted by Molybdenum-Rhenium (MoRe) electrodes.
 Bilayer graphene is obtained via the standard exfoliation method.
-It is then encapsulated in hexagonal Boron-Nitride using the dry transfer method [13].
+It is then encapsulated in hexagonal Boron-Nitride using the dry transfer method 14.
 MoRe of $80$ nm thickness is deposited via DC magnetron sputtering.
@@
 The width of the junctions is $4~{}\mu$m.
-The device is cooled in a Leiden cryogenics dilution refrigerator operated at temperatures above $1$ K, and the measurements were performed using the standard four-probe lock-in method.
-A gate voltage $V_{G}$ is applied to the $Si$ substrate with the oxide layer acting as a dielectric, which allows modulation of the carrier density.[14, 15, 4, 8, 16, 17].
-Figure 1: (a) Differential resistance versus gate voltage ($V_{G}$) and bias current $I_{Bias}$ taken at $T=1.37$ K.
+The device is cooled in a Leiden cryogenics dilution refrigerator operated at temperatures above $1$ K, and measurements were performed using the standard four-probe lock-in method.
+A gate voltage $V_{G}$ is applied to the $Si$ substrate with the oxide layer acting as a dielectric, which allows modulation of the carrier density.
+15, 16, 5, 2, 17, 6.
+Figure 1: (a) Differential resistance ($dV/dI$) versus gate voltage ($V_{G}$) and bias current $I_{Bias}$ taken at $T=1.37$ K.
 The black region around zero bias corresponds to the superconducting state.
-$I_{Bias}$ is swept from negative to positive.
+$I_{Bias}$ is swept up (from negative to positive).
 Thus, the transition at negative bias corresponds to the re-trapping current $I_{R}$, while the transition at positive bias is the switching current $I_{C}$.
-(b) Vertical line cut of the resistance map taken at $V_{G}=5$ V, $T=1.37$ K, showing device’s differential resistance versus bias current.
+(b) Vertical line cut of the resistance map taken at $V_{G}=15$ V, $T=1.37$ K, showing device’s $dV/dI$ versus bias current.
+Blue line corresponds to $I_{Bias}$ swept up, with red line swept down (positive to negative).
+Figure.
+1(a) displays the differential resistance ($dV/dI$) map of the $400$ nm junction at $T=1.37$ K; we see zero resistance (black region) across all applied $V_{G}$ indicating the presence of supercurrent.
+As the bias current $I_{bias}$ is swept from negative to positive values, the junction first reaches its superconducting state at a value $|I_{bias}|=I_{R}$, known as the re-trapping current.
+Then, as $|I_{bias}|$ is increased to higher positive values, the junction transitions to the normal state at $|I_{bias}|=I_{S}$, known as the switching current.
+Figure.
+1(a) shows that the junction can sustain a larger region of critical current as we modulate the carrier density to higher values via $V_{G}$.
 Fig.
-1(a) displays the differential resistance ($dV/dI$) map of the $400$ nm junction at $T=1.37$ K; we see zero resistance (black region) across all applied $V_{G}$ indicating the presence of supercurrent.
-As the bias current $I_{bias}$ is swept from negative to positive values, the junction first reaches its superconducting state at a value $|I_{bias}|=I_{R}$, known as re-trapping current.
-Then, as $|I_{bias}|$ is increased to higher positive values, the junction transitions to the normal state at $|I_{bias}|=I_{S}$, known as switching current.
-Fig.
-1(a) shows that the junction can sustain larger region of critical current as we modulate the carrier density to higher values via $V_{G}$.
-Fig.
-1(b) displays a line plot extracted from the $dV/dI$ map which shows hysteresis in $I_{R}$ and $I_{S}$.
-This is a commonly observed phenomenon in underdamped junctions [18, 14], or can also be attributed self-heating [15, 19, 16].
-The critical current $I_{C}$ of the junction is approximately $I_{S}$ as found in switching statistics measurements, discussed in previous publications [8, 20, 21, 22].
-Extracting the critical current $I_{C}$ from the differential maps for different temperatures, we can see that $I_{C}$ falls exponentially with inverse $T$ (Fig.
-2c) We also extract conductance of the junction in the normal regime ($I_{Bias}\gg I_{C}$).
-Fig.
-2(b) shows this conductance ($G$) for $400$ nm junction device.
-We find that conductance $G$ scales as the square-root of $V_{G}$ as seen from the fit (blue curve).
-However, due to large contact resistance ($R_{C}$) of the device, the measured conductance $G$ is suppressed compared to the ballistic limit expectation.
-Therefore, to demonstrate the ballistic nature of the device, we present normal resistances ($R_{N}$) of junctions of length $500$ nm, $600$ nm, and $700$ nm (Fig.
-2(c) inset).
-Taking into account $R_{C}$, we plot extracted values versus $V_{G}$ from the Dirac point.
-The inset plot shows that the $R_{N}-R_{C}$ values are independent of junction length, which demonstrates the ballistic nature of the devices.
+1(b) displays line traces extracted from the $dV/dI$ map which shows hysteresis in $I_{R}$ and $I_{S}$.
+This is a commonly observed phenomenon in underdamped junctions 18, 15, or can also be attributed to self-heating 16, 19, 17.
+The measured switching current $I_{S}$ is slightly suppressed compared to the junction’s “true” critical current $I_{C}$.
+However, previous measurements on the statistical distribution of $I_{S}$ in similar graphene devices found that $I_{S}$ is suppressed from $I_{C}$ by no more than $10\%$ for critical currents up to a few $\mu A$ 2, 20, 21, 22.
+Extracting the critical current $I_{C}$ from the differential maps for different temperatures, we can see that $I_{C}$ falls exponentially with inverse $T$ (Figure.
+2c) We also extract the conductance of the junction in the normal regime ($I_{Bias}\gg I_{C}$).
+Figure.
+2(b) shows this conductance ($G$) for the $400$ nm junction device.
+Due to the significant contact resistance ($R_{C}$) of the device, the measured conductance $G$ is uniformly suppressed compared to the ballistic limit expectation.
+However, when accounting for $R_{C}$ within the fit, we find that the conductance $G$ scales as the square-root (as opposed to linearly) of $V_{G}$ (blue curve of Figure 2(b)).
+This is consistent with ballistic transport23, 2.
+To further demonstrate the ballistic nature of the device, we present normal resistances ($R_{N}$) of junctions of length $500$ nm, $600$ nm, and $700$ nm with the fitted, constant contact resistance $R_{C}$ subtracted (Figure.
+2(b) inset).
+The inset plot shows that the values of $R_{N}-R_{C}$ are independent of the junction length, demonstrating the ballistic nature of the devices.
 Figure 2: (a) Device picture.
@@
 (b) The ballistic conductance vs Gate voltage for $L=400$ nm junction.
+The blue curve corresponds to the fit for ballistic devices, with an addition of a contact resistance.
 The inset shows junction resistance minus the parasitic contact resistance plotted against gate voltage from the Dirac point for all our devices.
@@
 The plots show $V_{G}$ dependence of $I_{C}$: the gray lines show that the slope of the curve for the lowest plotted gate $V_{G}=8$ V, is smaller than the slope of the highest plotted gate $V_{G}=21$ V.
-To extract $\delta E$ of the junction, we go to the discussion of $I_{C}$ vs temperature trends in Fig.
+To extract $\delta E$ of the junction, we go to the discussion of $I_{C}$ vs.
+the temperature trends in Figure.
 2(c).
-Here the y-axis is plotted in log scale.
-From the slope of curves $\mathrm{Log}(I_{C})=-(k_{B}/\delta E)T$ for each gate, one can extract $\delta E$ versus $V_{G}$ (plotted in Fig.
+Here, the y-axis is plotted in logarithmic scale.
+From the slope of the curves $\mathrm{Log}(I_{C})=-(k_{B}/\delta E)T$ for each gate, one can extract $\delta E$ versus $V_{G}$ (plotted in Figure.
 3a).
-Unlike for the case of monolayer graphene, a clear dependence on $V_{G}$ is seen.
-The energy $\delta E$ scales linearly with Fermi Velocity $v_{F}$ (Fig.
-3b).
+Unlike for the case of monolayer graphene, a clear dependence on $V_{G}$ is seen (The observed trend further supports the view that our devices operate in the long ballistic regime.
+Diffusive Josephson junctions are governed by the Thouless energy $E_{Th}\propto 1/[(R_{N}-R_{C})\sqrt{V_{G}-V_{D}}]$ 22, 24 which does not match the trend with respect to $V_{G}$ seen in Figure.
+3(a) ).
+The energy $\delta E$ scales linearly with the Fermi velocity $v_{F}$ (Figure.
+3(b) ).
 Note that calculating $v_{F}$ from $\delta E$ for junctions in the intermediate regime requires knowledge of the superconducting coherence length $\xi$.
-In our case $\xi$ is obtained from the fit discussed below.
+In the fit discussed below, we use $\xi$’s dependence in $v_{F}$.
 We now compare the experimentally obtained $\delta E$ (and $v_{F}$) to the theoretical expectation.
-Assuming a quadratic dispersion relation $\mathcal{E}_{F}=\frac{\hbar^{2}k^{2}_{F}}{2m^{*}}$, it follows that the expression for the Fermi velocity is: $v_{F}=\sqrt{\frac{2E_{F}}{m^{*}}}$ [23, 24, 25].
-The Fermi Energy $\mathcal{E}_{F}$ for bilayer graphene scales as: $\mathcal{E}_{F}=\frac{\hbar^{2}\pi|n|}{2m^{*}}$.
-With $m^{*}$ being the effective mass of electrons in graphene.
+With the dispersion relation for bilayer graphene written as : $\mathcal{E}=\frac{1}{2}\gamma_{1}(\sqrt{1+2\hbar^{2}k^{2}/\gamma_{1}^{2}m^{*}}-1)$, we get the expression for the Fermi velocity : $v_{F}=\sqrt{\frac{2\mathcal{E}_{F}\gamma_{1}(\mathcal{E}_{F}+\gamma_{1})}{(2\mathcal{E}_{F}+\gamma_{1})^{2}m^{*}}}$ 25, 26, 27.
+Here, $\gamma_{1}=0.39~{}eV$ a parameter describing the interlayer coupling 25, $k$ is the momentum wavevector, $m^{*}$ is the effective mass of electrons.
+Moreover, the Fermi energy $\mathcal{E}_{F}$ for bilayer graphene scales as: $\mathcal{E}_{F}=\frac{\hbar^{2}\pi|n|}{2m^{*}}$.
 The carrier concentration $n$, controlled by the applied gate voltage $V_{G}$, is given by $n=\frac{V_{G}-V_{D}}{e}C_{Total}$ with $V_{D}$ as the gate voltage at the Dirac point.
 The total capacitance $C_{Total}$ is a combination of quantum capacitance $C_{q}$ and gate oxide capacitance $C_{ox}$: $C_{Total}=\left[\frac{1}{C_{ox}}+\frac{1}{C{q}}\right]^{-1}$.
-The quantum capacitance $C_{q}$ for bilayer graphene is determined by $C_{q}=\frac{2e^{2}m^{*}}{\pi\hbar^{2}}$.
-The gate oxide capacitance per unit area is $C_{ox}=\frac{\epsilon_{0}\epsilon_{r}}{d}$, where $\epsilon_{0}$ is vacuum permittivity, $\epsilon_{r}$ is the relative permittivity of the oxide, and $d$ is oxide layer thickness.
+The quantum capacitance $C_{q}$ for bilayer graphene is determined by $C_{q}=\frac{2e^{2}m^{*}}{\pi\hbar^{2}}$, where $e$ is the electron charge.
+The gate oxide capacitance per unit area is $C_{ox}=\frac{\epsilon_{0}\epsilon_{r}}{d}$, where $\epsilon_{0}$ is the vacuum permittivity, $\epsilon_{r}$ is the relative permittivity of the oxide, and $d$ is the thickness of the oxide layer.
 For silicon oxide gate with $d=300$ nm we get $C_{ox}\approx 115\mu F/m^{2}$.
 Thus, the full expression for the Fermi velocity $v_{F}$ is:
-$v_{F}=\hbar\sqrt{\frac{2\pi\epsilon_{0}\epsilon_{r}e|V_{G}-V_{d}|}{m^{*}(2de^{2}m^{*}+\pi\epsilon_{0}\epsilon_{r}\hbar^{2})}}$ (1)
-Note that the effective mass $m*$ typically ranges from $0.024~{}m_{e}$ to $0.058~{}m_{e}$ for $1*10^{12}\sim 4*10^{12}$ carriers/$cm^{2}$ [26], where $m_{e}$ is the electron rest mass.
-Moreover, since $m*$ has a carrier concentration dependence, we assume a linear shift of $m*$ with the gate voltage as: $m^{*}=m_{i}+dm(V_{G}-V_{d})$ [26].
+$v_{F}=\hbar\sqrt{\frac{2\pi e\epsilon_{0}\epsilon_{r}\gamma_{1}(V_{G}-V_{D})(2de^{2}\gamma_{1}m^{*}+\pi\epsilon_{0}\epsilon_{r}\hbar^{2}(e(V_{G}-V_{D})+\gamma_{1}))}{m^{*}(2de^{2}\gamma_{1}m^{*}+\pi\epsilon_{0}\epsilon_{r}\hbar^{2}(2e(V_{G}-V_{D})+\gamma_{1}))^{2}}}$ (1)
+Note that the effective mass $m^{*}$ typically ranges from $0.024~{}m_{e}$ to $0.058~{}m_{e}$ for $1*10^{12}\sim 4*10^{12}$ carriers/$cm^{2}$ 28, where $m_{e}$ is the electron rest mass.
 Experimental data provides us with the following: $\delta E(V_{G})=\frac{\hbar}{2\pi(L+\xi)}v_{F}$.
-To fit this data, the model is set as :$\delta E(V_{G})=\mathcal{F}(m_{i},dm,\xi,V_{d},d)$ where $m_{i},dm,\xi,V_{d},d$ are the fitting parameters, and $V_{G}$ is the independent variable.
+We also note that $\xi$ has a dependence on $v_{F}$ and the superconducting gap $\Delta$: $\xi=\hbar v_{F}/2\Delta$13.
+To fit $\delta E$, the model is set as: $\delta E(V_{G})=\mathcal{F}(m^{*},\Delta,V_{D},d)$ where $m^{*},\Delta,V_{D},d$ are the fitting parameters, and $V_{G}$ is the independent variable.
 (We use the as-designed length of the device $L$, and take $\epsilon_{r}=3.9$ for $SiO_{2}$.)
-The resulting fit of the data from the $400~{}nm$ junction for both $\delta E$ and $v_{F}$ is plotted as solid lines in Fig.
-3(a) and Fig.
-3(b) respectively.
-Moreover, taking the fitted $\xi$, we calculate the Fermi velocity $v_{F}$ for all other junctions on the same substrate.
-As seen from Fig.
-3(b), the calculated $v_{F}$ of all devices is in good agreement with the fit obtained from the $400$ nm junction.
-The fitted parameters are summarized in Table 1.
-All fall within the range of expected values, with $\xi$ being consistent with previously measured values for graphene/MoRe junctions.
 Parameter Fitted Value Expected Value
-$\xi$ $460.34$ nm $300\sim 500$ nm
-$d$ $317.71$ nm $300\sim 330$ nm
-$m_{i}$ $0.020~{}m_{e}$ $0.02\sim 0.06~{}m_{e}$
-$dm$ $0.0003~{}m_{e}$/V $0.0002\sim 0.0005~{}m_{e}$/V
-$V_{d}$ $-1.99$ V ±$\approx\pm 2$ V
+$\Delta$ $0.99$ meV $0.8\sim 1.2$ meV
+$d$ $323$ nm $280\sim 330$ nm
+$m^{*}$ $0.028~{}m_{e}$ $0.02\sim 0.06~{}m_{e}$
+$V_{D}$ $2.04$ V $\approx+2$ V
 Table 1: The fitting parameters used to match the measured $\delta E$, and consequently the Fermi velocity $v_{F}$, versus gate to the theoretical expectation described in Equation 1.
@@
 The expected gate dielectric thickness $d$ is estimated from the substrate specifications plus the bottom hBN thickness.
-The expected Dirac point voltage $V_{D}$ is ontained from the resistance map.
-The expectations for the superconducting coherence length $\xi$ and the effective mass $m_{i}$ are obtained from previous works[26, 4].
+The expected Dirac point voltage $V_{D}$ is obtained from the resistance map.
+The expectations for superconducting gap $\Delta$ and the effective mass $m_{i}$ are obtained from previous works28, 5.
+The resulting fits of the data from the $400~{}nm$ junction for $\delta E$ and $v_{F}$ are plotted as solid lines in Figure.
+3(a) and Figure.
+3(b) respectively.
+Moreover, taking the fitted parameters from Table 1, we calculate the Fermi velocity $v_{F}$ for the available data points of all other junctions on the same substrate.
+As seen from Figure.
+3(b), the calculated $v_{F}$ of all devices is in good agreement with the fit obtained from the $400$ nm junction (This is as expected for devices on the same substrate; as long as they have consistent parasitic doping and superconductor-graphene contact interface).
+The fitted parameters are summarized in Table 1.
+All fall within the range of expected values, with $\Delta$ being consistent with previously measured values for graphene/MoRe junctions2.
+Furthermore, using the values obtained from the model, we find that $v_{F}$ saturates to the value of $1.1*10^{6}$ m/s as $V_{G}$ tends to infinity.
 Figure 3: (a) Energy $\delta E$ extracted from the slope of log($I_{C}$) vs T plotted against the gate voltage $V_{G}$ from the Dirac point of the junction with $L=400$ nm.
 We see $\delta E$ dependence on the carrier density modulated via the gate voltage for the junction.
-(b) Fermi velocity ($v_{f}$) calculated from $\delta E$ using the device dimensions, and the superconducting coherence length $\xi$ obtained from the fit to theory.
+(b) Fermi velocity ($v_{f}$) calculated from $\delta E$ using the device dimensions, and parameters obtained from the fit to theory.
 The solid line represents the theoretical trend as fitted to the data for the $L=400$ nm junction.
-In addition, panel (b) shows calculated $v_{F}$ for the other junctions using the $\xi$ from the $L=400$ nm fit.
-In conclusion, we study the evolution of critical current with respect to gate in bilayer graphene Josephson Junctions (BGJJs).
+In addition, panel (b) shows calculated $v_{F}$ for the other junctions using parameters obtained from the $L=400$ nm fit.
+In conclusion, we study the evolution of the critical current with respect to the gate in bilayer graphene Josephson Junctions (BGJJs).
 Using the critical current-temperature relation expected for intermediate-to-long junctions, we extract the relevant energy scale $\delta E$ and find that it has a clear gate dependence.
-As $\delta E$ is proportional to the Fermi velocity $v_{F}$ in the bilayer graphene, we are able to match the observed gate dependence to the theoretical expectation.
-Our observation is contrasted with monolayer graphene JJs, that do not have a gate dependent $\delta E$.
+As $\delta E$ is proportional to the Fermi velocity $v_{F}$ in bilayer graphene, we are able to match the observed gate dependence to the theoretical expectation.
+Our observation is contrasted with monolayer graphene JJs, which do not have a gate-dependent $\delta E$.
 This result showcases the greater tunability of BGJJs, and offers additional avenues for device characterization.
+Although not observed here, it should be possible to engineer Josephson junctions that transition from the short to the intermediate/long ballistic regimes in-situ via gate voltage.
+The ability to tune ABS level spacing could have applications in self-calibrating sensors, or for matching resonance conditions in multi-terminal superconducting devices.
+L.R. acknowledges support from NSF (DMR-2005092 award) for contact depositon. G.F. acknowledges support from Duke University. I.V.B. and A.S. acknowledge the support from Texas A&\&&M University.